OpenStudy (anonymous):

Using this equation y = x2 + 4x + 1. List your ordered pairs from the table. Use parenthesis for each pair, such as (-1,2).

OpenStudy (anonymous):

(-5,6), (-4,1), (-3,-2),(-2,-3), (-1,-2), (0,1), (1,6)<---- my calculator graphed it

OpenStudy (anonymous):

You can just pick an x, plug it into the equation, and see what y value you get. That will give you one ordered pair. Do that a bunch of times to get a list. Here's an example: I pick x = 0 y = (0)^2 +4(0) + 1 = 0 +0 +1 = 1 so I get the ordered pair (0,1)